A λ=3 MM SPECTRAL LINE SURVEY OF SAGITTARIUS B2(N-LMH) WITH THE BIMA ARRAY AND NRAO 12 METER RADIO TELESCOPE

\underline{DOUGLAS N. FRIEDEL}, LEWIS E. SNYDER, Department of Astronomy, University of Illinois, 1002 W. Green St., Urbana, IL 61801; BARRY E. TURNER, National Radio Astronomy Observatory, Charlottesville, VA 22903; and ANTHONY J. REMIJAN, Department of Astronomy, University of Illinois, 1002 W. Green St., Urbana, IL 61801 (email to D.N.F.: friedel@astro.uiuc.edu).

We have initiated a 3 mm spectral line survey toward the hot molecular core Sagittarius B2(N-LMH). This is the first spectral line survey of the Sgr B2(N) region utilizing data from both an interferometer (BIMA Array) and a single-element radio telescope (NRAO 12 meter). In this survey, covering 3.6 GHz in bandwidth, we detected 237 lines (90 identified molecular transitions, 1 electronic transition, and 148 unidentified transitions). Furthermore, the data provide information that could not be obtained by either instrument alone.